# Classification of ancient noncollapsed flows in $\mathbb{R}^{4}$

### v1

@@ 1. Introduction
The most important ancient solutions are the noncollapsed ones [And12, SW09, HK17], i.e. ancient mean-convex flows $M_{t}$, such that at any $p\in M_{t}$ the inscribed radius is at least $\alpha/H(p,t)$ for some uniform $\alpha>0$. It is known since the work of White [Whi00, Whi03] (see also [Whi15, HH18]) that all blowup limits of mean-convex mean curvature flow are noncollapsed. In fact, it is expected that any blowup limit near any generic singularity is noncollapsed [Ilm03, CM12]. In particular, this is known in $\mathbb{R}^{3}$ thanks to the recent proofs of the mean-convex neighborhood conjecture [CHH22], the genericity conjecture [CCMS24a, CCS23, CCMS24b] and the multiplicity-one conjecture [BK24b].
Understanding all noncollapsed singularities thus amounts to classifying all ancient noncollapsed flows. In $\mathbb{R}^{3}$, this classification problem has been studied intensively over the last 15 years, initially under additional assumptions such as selfsimilarity [Wan11, MSHS15, Has15, BW17], and a definite answer has been obtained in [ADS19, ADS20, BC19]:
@@ 1.1. The classification program in $\mathbb{R}^{4}$
Second, assuming one somehow solved the first challenge, one wants to show that every noncompact solution with mixed convergence is either $\mathbb{R}\times$ 2d-ovals or belongs to the 1-parameter family of $\mathbb{Z}_{2}\times\mathrm{O}_{2}$-symmetric entire translators from [HIMW19]. In light of the above, the second major challenge is thus:
@@
In general, the question whether an ancient (or eternal) noncompact solution is selfsimilar, is very subtle, as pointed out first by White [Whi03, Conjecture 1] and Perelman [Per02, Remark 11.9]. There are a some prior instances in the literature where selfsimilarity has been established successfully, namely [BC19, BC21, CHH22, CHHW22, Bre20, BN23]. However, the method of proof in all these prior papers was to establish rotational symmetry, and then selfsimilarity only followed a posteriori by uniqueness of rotationally symmetric solutions. Also, if one drops the noncollapsing condition then there are (many) examples of eternal strictly convex solutions that are not selfsimilar, see e.g. [BLT22].
@@ 1.2. The differential neck theorem and its consequences
Motivated by the analysis of the fast convergence, one might hope in the case of mixed convergence that $u$ as a function of $y_{1}$ still behaves like $ay_{1}e^{\tau/2}$. The problem with this approach is that the leading term $(2-y_{2}^{2})/(\sqrt{8}|\tau|)$ is way much larger, and thus trying to detect an exponentially small correction term sounds like a rather hopeless endeavour. However, we can turn this into a more meaningful approach, by instead considering the derivative of the profile function with respect to $y_{1}$, which kills the leading term and thus gives some hope to be able to detect the (dauntingly small) correction term.
@@ 1.3. Outline of the proofs
The only unstable eigenfunction for $\mathcal{L}^{\prime}$ is the constant function $1$, and the only $\mathrm{O}_{2}$-invariant neutral eigenfunctions of $\mathcal{L}^{\prime}$ are the linear functions $y_{1}$ and $y_{2}$. Denote the corresponding projections by $p^{\prime}_{\pm}$ and $p_{0}^{\prime}$. Fixing a suitable graphical radius function $\rho$ and a suitable cutoff function $\eta$, we work with the localized slope function
@@
On the other hand, after some decay such as (14) is available, we can iteratively apply barriers with increasing eccentricity alternated with our anisotropic propagation of smallness estimate (to be discussed below), yielding even more elongated cylindrical regions, specifically cylindrical regions of the form [^3] In particular, the this gives a (nonsharp) lower bound for cap distance, namely $d(t)\geq|t|^{1/2}(\log|t|)^{(k+1)/2}$.
@@
Namely, the estimate says that if $|\nu_{1}|$ decays like $e^{p\tau}$ in the core region $\{r_{\alpha}\leq L\}$, then we can propagate this decay to the larger region $\{r_{\alpha}\leq|\tau|^{1/2}\}$ up to loosing a small exponent. In particular, unwinding the definition of $r_{\alpha}$, note that in the $y_{1}$-variable we only have the factor $e^{\frac{1-\beta}{2}\alpha y_{1}^{2}}$, which is harmless for $\alpha$ small, so our estimate works particularly well to propagate smallness along the $y_{1}$-direction. Ping-ponging our anisotropic propagation of smallness estimate (with decreasing $\alpha$) and the anisotropic barrier estimate from above, we can upgrade (14) to $L^{\infty}$-decay in a long central strip, specifically we can show that
@@
We also have another propagation of smallness estimate, which allows us to propagate in $y_{2}$-direction from the central strip to the entire level sets, specifically we show that
@@
where $f_{\alpha}$ denotes the renormalized distance in the $(x,t)$-variables, and apply the maximum principle to the function $\psi=\eta w/\log|t|$, where $\eta$ is a carefully chosen weight function. To prove the level set propagation estimate (23), we first introduce a suitable supersolution of the Jacobi equation (24), specifically
@@
Yet another issue is that the level set propagation of smallness estimate only works provided that we have control for the diameter of the level sets. Namely, it only works provided that we have the spectral coefficient estimate $\langle 2-y_{2}^{2},u\rangle_{\mathcal{H}}\geq c|\tau|^{-1}$ for some $c>0$. To address this, we reconsider the classical Merle-Zaag analysis, and introduce a new variant that we call Merle-Zaag switch dynamics. It turns out that there is some switch time $\tau_{s}=\tau_{s}(X)$, where the spectral coefficients $\vec{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3})$ with respect to the eigenfunctions $y_{1}^{2}-2$, $y_{2}^{2}-2$ and $2y_{1}y_{2}$ reach some definite size $\kappa_{0}>0$. Specifically, we have
(27) $-\alpha_{2}\geq\tfrac{1}{2}\kappa_{0}|\tau|^{-1}\quad\mathrm{for}\;\;\tau\leq\tau_{s}\qquad\mathrm{and}\qquad|\vec{\alpha}|\leq\kappa_{0}|\tau|^{-1}\quad\mathrm{for}\;\;\tau_{s}<\tau\leq\tau_{\ast}$
So for $\tau\leq\tau_{s}$ we have some definite inwards quadratic bending, and thus can apply the level set propagation of smallness estimate. On the other hand, for $\tau_{s}<\tau\leq\tau_{\ast}$ we get a graphical radius much larger than $\sqrt{2|\tau|}$, so instead of (11) we can get an estimate with much smaller error terms, and we are fine as well.
Let us now sketch how we can establish our differential neck theorem by combining the above ingredients. We can assume that $\tau\leq\tau_{s}$, since in the easy case $\tau_{s}<\tau\leq\tau_{\ast}$ there is no iteration needed. Using (22) and the level set propagation of smallness, we get decay for the error term from (13), specifically we get $\omega(\tau)\leq C|\tau|e^{\frac{1}{10}\tau}$. Feeding this back into (11), we can improve (14) to $\|w\|_{\mathcal{H}}\leq Ce^{\frac{3}{10}\tau}$, which in turn yields a better decay for $\omega(\tau)$, etc. Bootstrapping this argument for several rounds we eventually obtain enough decay, so that considering the evolution of $a_{+}(\tau)=\langle w(\tau),1\rangle_{\mathcal{H}}$ we can obtain a weak version of the differential neck theorem. Specifically, we can show that there is some $a=a(\mathcal{M})\in\mathbb{R}$, such that
@@
Finally, let us outline how we can establish the classification theorem using the differential neck theorem. By scaling we can assume $a(\mathcal{M})=1/\sqrt{2}$. There is a unique point $p_{t}=p_{t}^{-}\in M_{t}$, where $x_{1}$ is minimal. In case $M_{t}$ is compact, there is also a unique point $p_{t}^{+}$, where $x_{1}$ is maximal. Denote by $\mathcal{T}(h)$ the time when a cap arrives at level $x_{1}=h$. This gives rise to a unique space-time point $P_{h}=(p^{\ast}_{\mathcal{T}(h)},\mathcal{T}(h))$. If $Z(P_{h_{i}})$ is bounded along some sequences $h_{i}\to\pm\infty$, then it is quite easy to conclude. Roughly speaking, in the compact case around $P_{h_{i}}$ with $h_{i}\to+\infty$ the slope would have the wrong sign, and in the noncompact case via Hamilton’s Harnack inequality [Ham95] we can show that $H(P_{h})\equiv 1$, since for $h_{i}\to\pm\infty$ we have convergence to translators with $a=1/\sqrt{2}$, hence with speed $1$. Ruling out the potential scenario that $Z(P_{h_{i}})\to\infty$ is more subtle. In that scenario, the ratio between the bubble-sheet scale and the curvature scale would go to infinity, namely $(ZH)(P_{h_{i}})\to\infty$. Fortunately, there are also some helpful features, specifically rescaling by $H(P_{h_{i}})$ around $P_{h_{i}}$ we would have convergence to the 3d-bowl, and rescaling by $Z(P_{h_{i}})^{-1}$ we would have convergence to $\mathbb{R}\times$2d-oval. The latter in turn yields $Q(P_{h})\leq CZ(P_{h})$, where $Q$ denotes the quadratic scale, i.e. the scale from which the inwards quadratic bending in $y_{2}$-direction starts to holds. To get a grasp on the potential scenario of unbounded bubble-sheet scale we consider the eccentricity at level $h$,
@@
Using the differential neck theorem, and some further propagation of smallness arguments, we show the translator ratio $\Theta=-\nu_{1}/H$ is close to $1$ on the whole level set, whenever the eccentricity is large. Specifically, we prove that there are $Z_{\star}<\infty$ and $N=N(\varepsilon)<\infty$, such that if $Z(P_{h})\geq Z_{\star}$, then for every height $h$ and time $t\leq\mathcal{T}(h)$ satisfying $\varsigma(h,t)\geq N+\frac{7}{4}\log Z(P_{h})$ we have

### v2

@@ 1. Introduction
The most important ancient solutions are the noncollapsed ones [And12, SW09, HK17], i.e. ancient mean-convex flows $M_{t}$, such that at any $p\in M_{t}$ the inscribed and exterior radius is at least $\alpha/H(p,t)$ for some uniform $\alpha>0$. It is known since the work of White [Whi00, Whi03] (see also [Whi15, HH18]) that all blowup limits of mean-convex mean curvature flow are noncollapsed. In fact, it is expected that any blowup limit near any generic singularity is noncollapsed [Ilm03, CM12]. In particular, this is known in $\mathbb{R}^{3}$ thanks to the recent proofs of the mean-convex neighborhood conjecture [CHH22], the genericity conjecture [CCMS24a, CCS23, CCMS24b] and the multiplicity-one conjecture [BK24b].
Understanding all noncollapsed singularities thus amounts to classifying all ancient noncollapsed flows. In $\mathbb{R}^{3}$ this classification problem has been studied intensively over the last 15 years, initially under additional assumptions such as selfsimilarity [Wan11, MSHS15, Has15, BW17], and a definite answer has been obtained in [ADS19, ADS20, BC19]:
@@ 1.1. The classification program in $\mathbb{R}^{4}$
Second, assuming one somehow solved the first challenge, one wants to show that every noncompact solution with mixed convergence is either $\mathbb{R}\times$ 2d-oval or belongs to the 1-parameter family of $\mathbb{Z}_{2}\times\mathrm{O}_{2}$-symmetric entire translators from [HIMW19]. In light of the above, the second major challenge is thus:
@@
In general, the question whether an ancient (or eternal) noncompact solution is selfsimilar is very subtle, as pointed out first by White [Whi03, Conjecture 1] and Perelman [Per02, Remark 11.9]. There are some prior instances in the literature where selfsimilarity has been established successfully, namely [BC19, BC21, CHH22, CHHW22, Bre20, BN23]. However, the method of proof in all these prior papers was to establish rotational symmetry, and then selfsimilarity only followed a posteriori by uniqueness of rotationally symmetric solutions. Also, if one drops the noncollapsing condition then there are (many) examples of eternal strictly convex solutions that are not selfsimilar, see [BLT22].
@@ 1.2. The differential neck theorem and its consequences
Motivated by the analysis of the fast convergence, one might hope in the case of mixed convergence that $u$ as a function of $y_{1}$ still behaves like $ay_{1}e^{\tau/2}$. The problem with this approach is that the leading term $(2-y_{2}^{2})/(\sqrt{8}|\tau|)$ is way larger, and thus trying to detect an exponentially small correction term sounds like a rather hopeless endeavour. However, we can turn this into a more meaningful approach, by instead considering the derivative of the profile function with respect to $y_{1}$, which kills the leading term and thus gives some hope to be able to detect the (still dauntingly small) correction term.
@@ 1.3. Outline of the proofs
The only unstable eigenfunction for $\mathcal{L}^{\prime}$ is the constant function $1$, and the only $\mathrm{O}_{2}$-invariant neutral eigenfunctions of $\mathcal{L}^{\prime}$ are the linear functions $y_{1}$ and $y_{2}$. We denote the corresponding projections by $p^{\prime}_{\pm}$ and $p_{0}^{\prime}$. Fixing a suitable graphical radius function $\rho$ and a suitable cutoff function $\eta$, we work with the localized slope function
@@
On the other hand, after some decay such as (14) is available, we can iteratively apply barriers with increasing eccentricity alternated with our anisotropic propagation of smallness estimate (to be discussed below), yielding even more elongated cylindrical regions, specifically cylindrical regions of the form [^3] In particular, this gives a (nonsharp) lower bound for cap distance, namely $d(t)\geq|t|^{1/2}(\log|t|)^{(k+1)/2}$.
@@
Namely, the estimate says that if $|\nu_{1}|$ decays like $e^{p\tau}$ in the core region $\{r_{\alpha}\leq L\}$, then we can propagate this decay to the larger region $\{r_{\alpha}\leq|\tau|^{1/2}\}$ up to losing a small exponent. In particular, unwinding the definition of $r_{\alpha}$, note that in the $y_{1}$-variable we only have the factor $e^{\frac{1-\beta}{2}\alpha y_{1}^{2}}$, which is harmless for $\alpha$ small, so our estimate works particularly well to propagate smallness along the $y_{1}$-direction. Ping-ponging our anisotropic propagation of smallness estimate (with decreasing $\alpha$) and the anisotropic barrier estimate from above, we can upgrade (14) to $L^{\infty}$-decay in a long central strip, specifically we can show that
@@
We also have another propagation of smallness estimate, which allows us to propagate smallness in $y_{2}$-direction from the central strip to the entire level sets, specifically we show that
@@
where $f_{\alpha}$ denotes the squared anisotropic distance in the $(x,t)$-variables, and apply the maximum principle to the function $\psi=\eta w/\log|t|$, where $\eta$ is a carefully chosen weight function. To prove the level set propagation estimate (23), we first introduce a suitable supersolution of the Jacobi equation (24), specifically
@@
Yet another issue is that the level set propagation of smallness estimate only works provided that we have control for the diameter of the level sets. Namely, it only works provided that we have the spectral coefficient estimate $\langle 2-y_{2}^{2},u\rangle_{\mathcal{H}}\geq c|\tau|^{-1}$ for some $c>0$. To address this, we reconsider the classical Merle-Zaag analysis, and introduce a new variant that we call Merle-Zaag switch dynamics. It turns out that there is some switch time $\tau_{s}(X)$, where the spectral coefficients $\vec{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3})$ with respect to the eigenfunctions $y_{1}^{2}-2$, $y_{2}^{2}-2$ and $2y_{1}y_{2}$ reach some definite size $\kappa_{0}>0$. Specifically, we have
(27) $-\alpha_{2}\geq\tfrac{1}{2}\kappa_{0}|\tau|^{-1}\quad\mathrm{for}\;\;\tau\leq\tau_{s}(X)\qquad\mathrm{and}\qquad|\vec{\alpha}|\leq\kappa_{0}|\tau|^{-1}\quad\mathrm{for}\;\;\tau_{s}(X)<\tau\leq\tau_{\ast}$
So for $\tau\leq\tau_{s}(X)$ we have some definite inwards quadratic bending, and thus can apply the level set propagation of smallness estimate. On the other hand, for $\tau_{s}(X)<\tau\leq\tau_{\ast}$ we get a graphical radius much larger than $\sqrt{2|\tau|}$, so instead of (11) we can get an estimate with much smaller error terms, and we are fine as well.
Let us now sketch how we can establish our differential neck theorem by combining the above ingredients. We can assume that $\tau\leq\tau_{s}(X)$, since in the easy case $\tau_{s}(X)<\tau\leq\tau_{\ast}$ there is no iteration needed. Using (22) and the level set propagation of smallness, we get decay for the error term from (13), specifically we get $\omega(\tau)\leq C|\tau|e^{\frac{1}{10}\tau}$. Feeding this back into (11), we can improve (14) to $\|w\|_{\mathcal{H}}\leq Ce^{\frac{3}{10}\tau}$, which in turn yields a better decay for $\omega(\tau)$, etc. Bootstrapping this argument for several rounds we eventually obtain enough decay, so that considering the evolution of $a_{+}(\tau)=\langle w(\tau),1\rangle_{\mathcal{H}}$ we can obtain a weak version of the differential neck theorem. Specifically, we can show that there is some $a=a(\mathcal{M})\in\mathbb{R}$, such that
@@
Finally, let us outline how we can establish the classification theorem using the differential neck theorem. By scaling we can assume $a(\mathcal{M})=1/\sqrt{2}$. There is a unique point $p_{t}=p_{t}^{-}\in M_{t}$, where $x_{1}$ is minimal. In case $M_{t}$ is compact, there is also a unique point $p_{t}^{+}$, where $x_{1}$ is maximal. Denote by $\mathcal{T}(h)$ the time when a cap arrives at level $x_{1}=h$. This gives rise to a unique space-time point $P_{h}=(p^{\ast}_{\mathcal{T}(h)},\mathcal{T}(h))$, which we call the cap at level $h$. If the bubble-sheet scale $Z(P_{h_{i}})$ is bounded along some sequences $h_{i}\to\pm\infty$, then it is quite easy to conclude. Roughly speaking, in the compact case around $P_{h_{i}}$ with $h_{i}\to+\infty$ the slope would have the wrong sign, and in the noncompact case via Hamilton’s Harnack inequality [Ham95] we can show that $H(P_{h})\equiv 1$, since for $h_{i}\to\pm\infty$ we have convergence to translators with differential neck constant $a=1/\sqrt{2}$, hence with speed $1$. Ruling out the potential scenario that $Z(P_{h_{i}})\to\infty$ is more subtle. In that scenario, the ratio between the bubble-sheet scale and the curvature scale would go to infinity, namely $(ZH)(P_{h_{i}})\to\infty$. Fortunately, there are also some helpful features, specifically rescaling by $H(P_{h_{i}})$ around $P_{h_{i}}$ we would have convergence to the 3d-bowl, and rescaling by $Z(P_{h_{i}})^{-1}$ we would have convergence to $\mathbb{R}\times$2d-oval. The latter in turn yields $Q(P_{h})\leq CZ(P_{h})$, where $Q$ denotes the quadratic scale, i.e. the scale from which the inwards quadratic bending in $y_{2}$-direction starts to holds. To get a grasp on the potential scenario of unbounded bubble-sheet scale we consider the eccentricity at level $h$,
@@
Using the differential neck theorem, and some further propagation of smallness arguments, we show the translator ratio $\Theta=-\nu_{1}/H$ is close to $1$ on the whole level set whenever the eccentricity is large. Specifically, we prove that there are $Z_{\star}<\infty$ and $N=N(\varepsilon)<\infty$, such that if $Z(P_{h})\geq Z_{\star}$, then for every height $h$ and time $t\leq\mathcal{T}(h)$ satisfying $\varsigma(h,t)\geq N+\frac{7}{4}\log Z(P_{h})$ we have